Circular No. 8443

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA PUPPIS 2004 S. Nakano, Sumoto, Japan, reports the independent discoveries of an

apparent nova by Akihiko Tago (Tsuyama, Okayama-ken, Japan; at mag 7.6 on two T-Max 400 films taken with a Pentax 67 camera on Nov. 20.672 UT), and Yukio Sakurai (Mito, Ibaragi-ken, Japan; two CCD frames taken on Nov. 20.812 using a Fuji ‘Fine Pix S2’ camera + Nikon 180-mm f/2.8 lens). Tago gave the variable’s position as α = 7^{h}42^{m}52^{s}, δ = −27^{o}06′35′′ (equinox 2000.0), adding that nothing was visible at this location on his films from Nov. 16.76 (limiting mag 11.5). Nakano measured position end figures 41^{m}53^{s}.76, 36′′.9 from Sakurai’s image, which shows the variable at mag 7.4; Sakurai notes that nothing was visible to mag 13.6 at this location on Nov. 12 and 16. Better-scale unfiltered CCD images yield position end figures 41^{m}53^{s}.56, 38′′.3 by R. Kushida (Yatsugatake South Base Observatory, 0.40-m Schmidt-Cassegrain f/10 reflector; mag 6.9 on Nov. 21.625) and 41^{m}53^{s}.59, 38′′.1 by K. Kadota (Ageo, Japan, 0.25-m f/5.0 reflector; mag 7.0 on Nov. 21.701). Kadota notes the presence of a star near this variable on Digitized Sky Survey plates taken in 1980, 1992, and 1995, with a USNO-A2.0 star having position end figures 41^{m}53^{s}.61, 39′′.1 (red mag 17.7, blue mag 18.2).

K. Ayani, Bisei Astronomical Observatory (BAO), reports that lowdispersion spectra (range 463–690 nm; resolution 0.54 nm at Hα) of the apparent nova, obtained on Nov. 21.75 at the BAO 1.01-m telescope, shows Hα and Hβ lines with P-Cyg profiles. Preliminary analysis of the Hα emission (without decomposition of the absorption component) yields FWHM = 650 km/s and equivalent width 0.6 nm, the minimum of the Hα P-Cyg absorption being blueshifted by 860 km/s with respect to the emission peak. Absorption of Fe II (multiplet 42) lines are clearly seen.

SUPERNOVAE 2004gd AND 2004ge Further to IAUC 8431, M. Moore, K. Shimasaki, and W. Li report the

LOSS discovery of two apparent supernovae on unfiltered KAIT images:

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004gd Nov. 6.45 7^{h}09^{m}11.^{s}71 +20^{o}36′10′′.6 17.3 3′′.7 W, 0′′.8 N
2004ge Nov. 17.46 6 50 00.13 +25 38 02.0 18.3 6′′.2 E, 1′′.4 S

Additional photometry: SN 2004gd in NGC 2341, Oct. 22.46 UT, [19.5; Nov. 14.44, 17.2; 17.45, 17.2; 20.44, 17.1. SN 2004ge in UGC 3555, Nov. 14.45 UT, [19.5; 20.44, 17.8.

2004 November 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green